# Influence of temperature-dependent density inhomogeneity on the stability of atmospheric stratified fluids

### v2

@@ I Introduction
Continued global warming and climate change deserve great interest from scientists studying problems of atmospheric circulation dynamics, e.g., problems describing free thermal convection Busse (1977). Such motions, which arise in the gravity field at spatial inhomogeneity of density caused by temperature inhomogeneity were intensively discussed. Acheson and Hide (Acheson and Hide, 1973) noted that differential heating produces temperature variations from place to place, and owing to thermal expansion, they give rise to density stratification, causing so-called thermal instabilities. In Tur and Yanovsky (2013), the effect of a small-scale helical driving force on fluid with a stable vertical temperature gradient, with a small Reynolds number, under the action of gravity is considered. The authors showed large-scale vortex instability in the fluid despite its stable stratification. At a nonlinear stage, this instability gives many stationary spiral vortex structures. Among these structures, there is a stationary helical soliton and a kink of the new type. Later in Tur et al. (2013), authors considered stratified in gravity field rotating flow and found a new large-scale instability. They obtained nonlinear equations for the instability and gave a detailed study of the linear stage of the instability and the conditions of its appearance. They noted that instability is possible in the case of both stable and unstable vertical temperature stratifications. The authors in Kopp et al. (2021) reported a new type of large-scale instability generated by the unstable vertical temperature gradient (heated from below) and small-scale force with zero helicity in an inclined rotating fluid with solutions in localized vortex kink-like structures. Such investigations dealt with the propagation problems of internal gravity waves Sutherland (2010). In Ref. Yiğit and Medvedev (2015), the authors discussed vertical coupling in the atmosphere and ionosphere system by internal waves generated in the lower atmosphere. In addition, they also discussed the progress in the sudden stratospheric warming and upper atmospheric circulation due to wave-induced vertical coupling between the lower and upper atmosphere.
@@
Recently, Kaladze and Misra Kaladze and Misra (2023) studied the influences of the temperature and density gradients due to thermal expansion on the stability of vertically stratified fluids in the Earth’s atmosphere. The authors initiated an investigation in the atmospheric layer ($500<z<$ 50 km heights) with negative and positive vertical temperature gradients in the tropospheric ($150<z<$ 15 km) and stratospheric $5015<z<$ 50 km) regions separately. They showed that such influence could lead to instability in stratified incompressible fluids. They also obtained the Brunt-Väisälä frequency modified by the thermal expansion coefficient and the critical value of the expansion coefficient for which the instability of internal gravity waves occurs. In their approach, they ignored the dependence of the background density of stratified fluids on the variation of the background fluid temperature. However, spatial variations of fluid temperatures in the atmosphere can cause density variations owing to the thermal expansion Acheson and Hide (1973). Such temperature dependence can significantly influence the spatio-temporal evolution of perturbed flows. Thus, it becomes necessary to take into account such effects in the dynamics of stratified atmospheric fluids.
@@
In this work, we consider such density inhomogeneity effect due to temperature variation and the vertical temperature gradient with an arbitrary sign, which allows the expansion of atmospheric layers with heights ranging from 00 to $50$ km and to consider arbitrary length scales of density and temperature inhomogeneities. In this way, we modify and advance the work by Kaladze and Misra Kaladze and Misra (2023). We show that the Brunt-Väisälä frequency associated with the internal gravity waves (IGWs) gets modified by this effect, and we obtain new instability conditions for IGWs. Consequently, the Rayleigh-Bénard convective instability and the modified instability growth rate occur.
@@ II Modeling of density inhomogeneity due to thermal expansion
Typically, the thermal expansion coefficient, $\beta$ is small $(\ll 1)$, and it decreases with increasing the temperature Kaladze and Misra (2023). So, the condition (4) gives
@@ III Basic Equations
Equations (12), (13), (17), and (18) are the desired set of equations for the dynamics of atmospheric stratified incompressible fluid flows. Comparing these equations with those in Kaladze and Misra (2023), we note that not only the effect of fluid viscosity is considered in addition to the present work, the temperature (unperturbed) gradient $(C_{0})$ appears to be constant in contrast to the work Kaladze and Misra (2023).
@@ IV Brunt-Väisälä frequency and instability condition
Equation (29) has the same form as for ordinary internal gravity waves (Stenflo, 1987; Kaladze et al., 2022) in absence of the effects of thermal expansion. It is important to note that the choice of the mean fluid density (background) is crucial for the spatio-temporal evolution of the perturbed fluid velocity $w$. Since the temperature variation in the Earth’s atmosphere can cause the density variation due to thermal expansion, we model the background density $\overline{\rho}(z)\equiv\overline{\rho}\left(\overline{T}(z),\overline{p}(z)\right)$ as Acheson and Hide (1973)
@@
where $\rho_{0}\equiv\rho_{0}(z)$, ${\cal T}\equiv{\cal T}(z)=\overline{T}(z)-T_{0}$. Comparing the first form of Eq. (32) with Eq. (21) of Ref. (Kaladze and Misra, 2023), we find that an additional factor $1-\beta{\cal T}(z)$ appears in Eq. (32) before the square brackets due to the modeling of the background density with thermal expansion effects [cf. Eq. (31)]. Furthermore, in contrast to Ref. (Kaladze and Misra, 2023), the temperature gradient $C_{0}\equiv d\overline{T}(z)/dz$ is no longer varying but a constant [cf. Eq. (16)].
In what follows, we obtain the instability conditions from Eq. (32), and note that $d\rho_{0}/dz<0$ holds in the atmospheric heights $500<z<$ 50 km Kaladze and Misra (2023) and $\beta{\cal T}<1$ holds good as per the assumption (9). Furthermore, ${d{\cal T}}/{dz}<0$ in the tropospheric heights $150<z<$ 15 km and ${d{\cal T}}/{dz}>0$ in the stratospheric heights $5015<z<$ 50 km Kaladze and Misra (2023). So, in the tropospheric layer, the necessary conditions for the instability at which $N^{2}<0$ are $\beta{\cal T}<1$ and ${d{\cal T}}/{dz}<0$. On the other hand, since the conditions ${d{\cal T}}/{dz}>0$ and $\beta{\cal T}<1$ are likely to hold in the stratosphere, the fluid is said to be stable there for $N^{2}>0$. The sufficient condition of instability gives
@@
Since as said before, $d\rho_{0}/dz<0$ in the atmospheric region of heights $500<z<$ 50 km and $\beta{\cal T}<1$ as per the assumption (9), the condition (33) can only be satisfied when ${d{\cal T}}/{dz}<0$. Typically, the length scale of the temperature inhomogeneity is larger than that of the density inhomogeneity Kaladze and Misra (2023), i.e.,
@@
The condition (35) is satisfied only when $0.5\lesssim\beta{\cal T}\lesssim 1$, which is rather a harsh one in the atmospheric regions. Thus, we may conclude that incompressible stratified fluid flows may be unstable in the tropospheric layer $150<z<$ 15 km but stable in the other region, i.e., $5015<z<$ 50 km. For typical values of the atmospheric fluid density $\rho_{0}(z)$ and temperature $\overline{T}$, and their gradients, readers are referred to the work of Kaladze and Misra Kaladze and Misra (2023). Using the data as in Table 1 of Ref. Kaladze and Misra (2023), it is seen from Fig. 1 that for a fixed temperature gradient, a critical value of the thermal expansion coefficient $\beta$ exists close to or above $(\beta\gtrsim 0.004/\mathrm{K})$ which the instability of stratified fluid may occur in the tropospheric layer. Such a critical value is in agreement with the previous work Kaladze and Misra (2023).
Figure 1: The squared Brunt-Väisälä frequency ($N^{2}$) is plotted against the height ($z$) for different values of the thermal expansion coefficient $\beta$ as in the legend. The data used are as in Table 1 of Ref. Kaladze and Misra (2023). It is seen that for a fixed temperature gradient $dT_{0}/dz~{}(\mathrm{K/m})$, the instability occurs $(N^{2}<0)$ for $\beta\gtrsim 0.004/\mathrm{K}$, otherwise the stratified fluid is said to be stable $(N^{2}>0)$.
@@ V Rayleigh-Bénard convective instability
and $K_{N}=\left(k^{4}+N^{4}k_{z}^{2}/g^{2}\right)^{1/4}$. From Eqs. (50) and (51), we note that for $N^{2}<0$ and $k_{z}>0$, we have the instability $(\gamma>0)$ and a real wave mode propagating vertically upwards in the atmosphere with the frequency $\omega_{r}~{}(>0)$. Typically, for a particular atmospheric model, $|N^{2}|\sim 10^{-4}-10^{-3}/\rm{s}^{2}$ and $g\sim 10~{}\rm{m/s^{2}}$ Kaladze and Misra (2023) so that $K_{N}\approx k$ for $k_{z}\ll 1$ such that $\omega_{r}\approx 0$. However, for moderate values of $|N^{2}|$ and $k_{z}$, such an approximation may not be valid. Thus, in this approximation and in absence of the effects of fluid viscosity and thermal diffusivity, the Rayleigh-Bénard convective flows with positive lapse rate $\Gamma$ can no longer propagate with a finite wave frequency $\omega_{r}$, but can have a finite growth rate of instability, given by,
@@
Clearly, the instability grows and the instability growth rate increases with increasing values of the thermal coefficient $\beta$ and the thermal lapse rate $\Gamma$ without any cut-off unless $k_{\perp}=0$. Since the thermal lapse rate remains a constant, the increasing values of $\beta$ may correspond to the region of low-temperatures with heights ranging from 00 to $15$ km (See, e.g., Table 1 of Ref. Kaladze and Misra (2023)).
@@ VI Conclusion
We have investigated the influence of the temperature-dependent density inhomogeneity due to thermal expansion on the stability of stratified fluids in the Earth’s lower atmosphere ($500<z<$ 50 km) owing to its importance in the spatio-temporal evolution of perturbed flows. We show that modeling of such temperature dependence introduces a new factor involving the thermal expansion coefficient in the Brunt-Väisälä frequency and thus modifies its expression and the instability condition compared to those proposed in Ref. Kaladze and Misra (2023). Furthermore, in contrast to Ref. Kaladze and Misra (2023), the instability of stratified flows can occur for $\beta{\cal T}<1$ and $d\overline{T}/dz<0$ only in the tropospheric region $(0<z<15)$ km. At the stratospheric region $(15<z<50)$ km, the fluid flow can be stable. It is also noted that the background temperature gradient, previously considered to vary, does not vary with atmospheric heights but remains a constant. We also consider the influence of the modified Brunt-Väisälä frequency, $N^{2}$ on the Rayleigh-Bénard convective flows and show that the negative temperature gradient for which the stratified flows associated with internal gravity waves become unstable can cause convective instability with an enhanced growth rate at a lower value of $N^{2}$. Such growth rates can be higher with increasing values of the thermal expansion coefficient and the thermal lapse rate, but can be significantly reduced by the effects of the thermal diffusion.
@@
To conclude, the density inhomogeneity-driven thermal instability of stratified fluids relating to the Brunt-Väisälä frequency could be helpful for the initiation of large-scale (may be comparable to the turbulence or vortex phenomena) instabilities through the momentum and energy transfer such as those associated with internal gravity waves and acoustic-gravity waves in the atmosphere Rogachevskii and Kleeorin (2024). Furthermore, due to the Rayleigh-Bénard convective instability, a more efficient energy transfer between two layers in the atmosphere can take place by the effects of the thermal expansion and the temperature gradient, leading to a transition from laminar to turbulent flows Zhao et al. (2024); Sharifi Ghazijahani and Cierpka (2024). However, such efficient energy transport may eventually tend to resist some external perturbations, and the fluid flows may return to the equilibrium state. However, the detailed discussion concerning it is beyond the scope of the present study.

### v3

@@ I Introduction
Continued global warming and climate change deserve great interest from scientists studying problems of atmospheric circulation dynamics, e.g., problems describing free thermal convection [1]. Such motions, which arise in the gravity field at spatial inhomogeneity of density caused by temperature inhomogeneity were intensively discussed. Acheson and Hide [2] noted that differential heating produces temperature variations from place to place, and owing to thermal expansion, they give rise to density stratification, causing so-called thermal instabilities. In [3], the effect of a small-scale helical driving force on fluid with a stable vertical temperature gradient, with a small Reynolds number, under the action of gravity is considered. The authors showed large-scale vortex instability in the fluid despite its stable stratification. At a nonlinear stage, this instability gives many stationary spiral vortex structures. Among these structures, there is a stationary helical soliton and a kink of the new type. Later in [4], authors considered stratified in gravity field rotating flow and found a new large-scale instability. They obtained nonlinear equations for the instability and gave a detailed study of the linear stage of the instability and the conditions of its appearance. They noted that instability is possible in the case of both stable and unstable vertical temperature stratifications. The authors in [5] reported a new type of large-scale instability generated by the unstable vertical temperature gradient (heated from below) and small-scale force with zero helicity in an inclined rotating fluid with solutions in localized vortex kink-like structures. Such investigations dealt with the propagation problems of internal gravity waves [6]. In Ref. [7], the authors discussed vertical coupling in the atmosphere and ionosphere system by internal waves generated in the lower atmosphere. In addition, they also discussed the progress in the sudden stratospheric warming and upper atmospheric circulation due to wave-induced vertical coupling between the lower and upper atmosphere.
@@
Recently, Kaladze and Misra [8] studied the influences of the temperature and density gradients due to thermal expansion on the stability of vertically stratified fluids in the Earth’s atmosphere. The authors initiated an investigation in the atmospheric layer ($500<z<$ 50 km heights) with negative and positive vertical temperature gradients in the tropospheric ($150<z<$ 15 km) and stratospheric $5015<z<$ 50 km) regions separately. They showed that such influence could lead to instability in stratified incompressible fluids. They also obtained the Brunt-Väisälä frequency modified by the thermal expansion coefficient and the critical value of the expansion coefficient for which the instability of internal gravity waves occurs. In their approach, they ignored the dependence of the background density of stratified fluids on the variation of the background fluid temperature. However, spatial variations of fluid temperatures in the atmosphere can cause density variations owing to the thermal expansion [2]. Such temperature dependence can significantly influence the spatio-temporal evolution of perturbed flows. Thus, it becomes necessary to take into account such effects in the dynamics of stratified atmospheric fluids.
@@
In this work, we consider such density inhomogeneity effect due to temperature variation and the vertical temperature gradient with an arbitrary sign, which allows the expansion of atmospheric layers with heights ranging from 00 to $50$ km and to consider arbitrary length scales of density and temperature inhomogeneities. In this way, we modify and advance the work by Kaladze and Misra [8]. We show that the Brunt-Väisälä frequency associated with the internal gravity waves (IGWs) gets modified by this effect, and we obtain new instability conditions for IGWs. Consequently, the Rayleigh-Bénard convective instability and the modified instability growth rate occur.
@@ II Modeling of density inhomogeneity due to thermal expansion
Typically, the thermal expansion coefficient, $\beta$ is small $(\ll 1)$, and it decreases with increasing the temperature [8]. So, the condition (4) gives
@@ III Basic Equations
Equations (12), (13), (17), and (18) are the desired set of equations for the dynamics of atmospheric stratified incompressible fluid flows. Comparing these equations with those in [8], we note that not only the effect of fluid viscosity is considered in addition to the present work, the temperature (unperturbed) gradient $(C_{0})$ appears to be constant in contrast to the work [8].
@@ IV Brunt-Väisälä frequency and instability condition
Equation (29) has the same form as for ordinary internal gravity waves [9, 10] in absence of the effects of thermal expansion. It is important to note that the choice of the mean fluid density (background) is crucial for the spatio-temporal evolution of the perturbed fluid velocity $w$. Since the temperature variation in the Earth’s atmosphere can cause the density variation due to thermal expansion, we model the background density $\overline{\rho}(z)\equiv\overline{\rho}\left(\overline{T}(z),\overline{p}(z)\right)$ as [2]
@@
where $\rho_{0}\equiv\rho_{0}(z)$, ${\cal T}\equiv{\cal T}(z)=\overline{T}(z)-T_{0}$. Comparing the first form of Eq. (32) with Eq. (21) of Ref. [8], we find that an additional factor $1-\beta{\cal T}(z)$ appears in Eq. (32) before the square brackets due to the modeling of the background density with thermal expansion effects [cf. Eq. (31)]. Furthermore, in contrast to Ref. [8], the temperature gradient $C_{0}\equiv d\overline{T}(z)/dz$ is no longer varying but a constant [cf. Eq. (16)].
In what follows, we obtain the instability conditions from Eq. (32), and note that $d\rho_{0}/dz<0$ holds in the atmospheric heights $500<z<$ 50 km [8] and $\beta{\cal T}<1$ holds good as per the assumption (9). Furthermore, ${d{\cal T}}/{dz}<0$ in the tropospheric heights $150<z<$ 15 km and ${d{\cal T}}/{dz}>0$ in the stratospheric heights $5015<z<$ 50 km [8]. So, in the tropospheric layer, the necessary conditions for the instability at which $N^{2}<0$ are $\beta{\cal T}<1$ and ${d{\cal T}}/{dz}<0$. On the other hand, since the conditions ${d{\cal T}}/{dz}>0$ and $\beta{\cal T}<1$ are likely to hold in the stratosphere, the fluid is said to be stable there for $N^{2}>0$. The sufficient condition of instability gives
@@
Since as said before, $d\rho_{0}/dz<0$ in the atmospheric region of heights $500<z<$ 50 km and $\beta{\cal T}<1$ as per the assumption (9), the condition (33) can only be satisfied when ${d{\cal T}}/{dz}<0$. Typically, the length scale of the temperature inhomogeneity is larger than that of the density inhomogeneity [8], i.e.,
@@
The condition (35) is satisfied only when $0.5\lesssim\beta{\cal T}\lesssim 1$, which is rather a harsh one in the atmospheric regions. Thus, we may conclude that incompressible stratified fluid flows may be unstable in the tropospheric layer $150<z<$ 15 km but stable in the other region, i.e., $5015<z<$ 50 km. For typical values of the atmospheric fluid density $\rho_{0}(z)$ and temperature $\overline{T}$, and their gradients, readers are referred to the work of Kaladze and Misra [8]. Using the data as in Table 1 of Ref. [8], it is seen from Fig. 1 that for a fixed temperature gradient, a critical value of the thermal expansion coefficient $\beta$ exists close to or above $(\beta\gtrsim 0.004/\mathrm{K})$ which the instability of stratified fluid may occur in the tropospheric layer. Such a critical value is in agreement with the previous work [8].
Figure 1: The squared Brunt-Väisälä frequency ($N^{2}$) is plotted against the height ($z$) for different values of the thermal expansion coefficient $\beta$ as in the legend. The data used are as in Table 1 of Ref. [8]. It is seen that for a fixed temperature gradient $dT_{0}/dz~{}(\mathrm{K/m})$, the instability occurs $(N^{2}<0)$ for $\beta\gtrsim 0.004/\mathrm{K}$, otherwise the stratified fluid is said to be stable $(N^{2}>0)$.
@@ V Rayleigh-Bénard convective instability
and $K_{N}=\left(k^{4}+N^{4}k_{z}^{2}/g^{2}\right)^{1/4}$. From Eqs. (50) and (51), we note that for $N^{2}<0$ and $k_{z}>0$, we have the instability $(\gamma>0)$ and a real wave mode propagating vertically upwards in the atmosphere with the frequency $\omega_{r}~{}(>0)$. Typically, for a particular atmospheric model, $|N^{2}|\sim 10^{-4}-10^{-3}/\rm{s}^{2}$ and $g\sim 10~{}\rm{m/s^{2}}$ [8] so that $K_{N}\approx k$ for $k_{z}\ll 1$ such that $\omega_{r}\approx 0$. However, for moderate values of $|N^{2}|$ and $k_{z}$, such an approximation may not be valid. Thus, in this approximation and in absence of the effects of fluid viscosity and thermal diffusivity, the Rayleigh-Bénard convective flows with positive lapse rate $\Gamma$ can no longer propagate with a finite wave frequency $\omega_{r}$, but can have a finite growth rate of instability, given by,
@@
Clearly, the instability grows and the instability growth rate increases with increasing values of the thermal coefficient $\beta$ and the thermal lapse rate $\Gamma$ without any cut-off unless $k_{\perp}=0$. Since the thermal lapse rate remains a constant, the increasing values of $\beta$ may correspond to the region of low-temperatures with heights ranging from 00 to $15$ km (See, e.g., Table 1 of Ref. [8]).
@@ VI Discussion and conclusion
We have investigated the influence of the temperature-dependent density inhomogeneity due to thermal expansion on the stability of stratified fluids in the Earth’s lower atmosphere ($500<z<$ 50 km) owing to its importance in the spatio-temporal evolution of perturbed flows. We show that modeling of such temperature dependence introduces a new factor involving the thermal expansion coefficient in the Brunt-Väisälä frequency and thus modifies its expression and the instability condition and regions compared to those proposed in Ref. [8]. Furthermore, in contrast to Ref. [8], the instability of stratified flows can occur for $\beta{\cal T}<1$ and $d\overline{T}/dz<0$ only in the tropospheric region $(0<z<15)$ km. Consequently, the wave-eigenmode and the instability growth rate in the troposphere will be modified, and so will the evolution and the dynamics of internal gravity waves, governed by Eq. (29), in the vertical stratified fluids. The detailed discussion along these lines is beyond the scope of the present investigation but a project of future work. In the stratospheric region $(15<z<50)$ km, the fluid flow can be stable. It is also noted that the background temperature gradient, previously considered to vary in Ref. [8], does not vary with atmospheric heights but remains a constant according to our assumption.
We also consider the influence of the modified Brunt-Väisälä frequency, $N^{2}$ on the Rayleigh-Bénard convective flows and show that the negative temperature gradient for which the stratified flows associated with internal gravity waves become unstable can cause convective instability with an enhanced growth rate at a lower value of $N^{2}$. Such growth rates can be higher with increasing values of the thermal expansion coefficient and the thermal lapse rate, but can be significantly reduced by the effects of the thermal diffusion.
Our basic model has some limitations. For example, in the present investigation, we have neglected the fluid density variation due to the pressure inhomogeneity compared to the temperature inhomogeneity and the Earth’s rotation. Inclusion of these effects could again modify the Brunt-Väisälä frequency, and so is the instability condition of stratified fluids. Furthermore, our modeling of the temperature-dependent density inhomogeneity and the absence of any heat source give rise to the constant gradient of the unperturbed temperature. However, the data in Ref. [8] suggests that the temperature gradient may not be precisely constant but may vary slightly with the Atmospheric height. We have also limited our discussion on the Rayleigh-Bénard convective instability to unbounded regions without estimating a critical Raleigh number beyond which the instability may occur. However, since the temperature can change sharply with the atmospheric height, the tropospheric region may be considered to be bounded between two heights with two different temperatures. Consideration of such bounded region together with the pressure inhomogeneity in the solution of Eq. (47) could significantly change the critical Raleigh number and the Nusselt number for critical modes and hence the instability growth rates. We have also noted the possibility of thermal waves having a real-wave frequency and an instability-growth rate. Thus, one can calculate the vertical average heat flux for such unstable thermal waves to estimate the Nusselt number for critical modes. Such unstable thermal waves at the nonlinear stage can open up several challenging areas to explore. However, further investigations in these lines are underway, and we will communicate them elsewhere. Furthermore, since we have limited our discussion, particularly to the modification of the Brunt-Väisälä frequency and hence the instability condition of stratified fluids, as well as the Rayleigh-Bénard convective instability of linear wave modes, the direct numerical simulation of Eqs. (12), (13), (17), and (18) have not been carried out but may be necessary for nonlinear evolution in our future studies.
The ongoing rise in global temperature significantly affects the Earth’s climate system, and its change can affect several atmospheric variables and the dynamics of atmospheric waves, e.g., internal gravity waves associated with them. The jump in temperature due to the absorption and release of heat by greenhouse gases causes the lower atmosphere (troposphere) to expand. Thus, predicting the instability of vertically stratified fluids due to thermal expansion and temperature gradients in the troposphere reported here could be crucial for realizing large-scale instability (may be larger than the scale of vortex or turbulence phenomena) through which the particle momentum and energy transfer can occur.
As stated before, the proper modeling of the density inhomogeneity due to thermal expansion modifies the previously known instability condition [8], i.e., the instability of stratified fluids may not occur in the entire region $500<z<$ 50 km but more specifically in the tropospheric region: $150<z<$ 15 km. Thus, the present theory may more precisely predict the consequences of thermal instability (e.g., the nonlinear evolution and the dynamics of internal gravity waves) due to climate changes in the tropospheric region than that presented in Ref. 8.
@@
To conclude, the density inhomogeneity-driven thermal instability of stratified fluids relating to the Brunt-Väisälä frequency could be helpful for the initiation of large-scale (may be comparable to the turbulence or vortex phenomena) instabilities through the momentum and energy transfer such as those associated with internal gravity waves and acoustic-gravity waves in the atmosphere [11]. Furthermore, due to the Rayleigh-Bénard convective instability, a more efficient energy transfer between two layers in the atmosphere can take place by the effects of the thermal expansion and the temperature gradient, leading to a transition from laminar to turbulent flows [12, 13]. However, such efficient energy transport may eventually tend to resist some external perturbations, and the fluid flows may return to the equilibrium state. However, the detailed discussion concerning it is beyond the scope of the present study.
Acknowledgments
The authors thank the anonymous Referees for their insightful comments, which improved the manuscript in its present form.